A man starts saving on 1st November, he save Shs 1 the first day,Shs.2 on second day Shs.4 the third day and so on doubling the amount everyday. If he managed to keep on saving until the end of month, how much would he saved?
Answer by fractalier(6550)

Thirty days in November...first day is 2^0 = 1, second day is 2^1 = 2...so that the 30th day is 2^29...thus the total turns out to be
2^30 - 1 = 1,073,741,823